# MobileVLM: A Vision-Language Model for Better Intra- and Inter-UI Understanding

### v1

@@ C.2 Stage2-CheckPoints Result
Figure 14: Stage2 Metrics Over Training Checkpoints.
@@ Appendix D Case Study
D.1 Grounding Task
Comparing Figure 18 and Figure 18, it can be observed that Qwen-VL performs well on the original object detection tasks. However, after undergoing Stage 1 training, the LLM experienced a decline in performance on these tasks. However, comparing Figures 18 and 18, it can be seen that the model trained with Stage 1 shows improvement on the Mobile3M Grounding task compared to the original Qwen-VL. In the selected examples, Qwen had significant errors, but Stage 1 provided a corrective effect across a broader range of examples. Considering that the data is suitable for training specialized models in the Mobile domain, the performance loss in the general domain is acceptable.
@@ D.1 Grounding Task
D.2 Pre-training Task Test Case
@@
2. Action Space Generation As shown in Figure 20, based on the Element List, we expect the model to further distinguish the type of each widget. Since scroll elements are not visible themselves, we manually generated four alternative actions for them: up, down, left, and right. Our action space does not support diagonal scrolling, as this often implies a drag action in practical operations. Therefore, the model actually only needs to distinguish between input fields and buttons. We also expect it to accurately ground these buttons.

### v2

@@ C.1 Stage1-CheckPoints Result
Figure 15: Stage2 Metrics Over Training Checkpoints.
@@ Appendix D Case Study
D.1 Illusrtration for Training Consistency
The consistency here mainly includes two aspects: (1) Both our pre-training data and fine-tuning domain data are sourced from the mobile domain. (2) The tasks during both the pre-training and fine-tuning stages are specifically designed for the mobile domain. Below, we will use examples to illustrate this issue.
Table 16: Comparison of Training Frameworks
Previous Training Framework Mobile3M Training Framework
Pre-training stage1 Grounding tasks on refCOCO. Element Grounding tasks on Mobile3M.
Data Example “Where is the girl petting the dog?” “Where is the input control containing Hangzhou in the screenshot?”
Pre-training stage2 VQA tasks on ChartQA. Action Prediction tasks on Mobile3M.
Data Example “What is the result of the addition calculation in the table?” “What action should I take to go from image one to image two?”
Fine-tuning Instruction navigation tasks on Auto-UI. Instruction navigation tasks on Mobile3M.
Data Example “How should I open the alarm clock?” “How should I open the alarm clock?”
D.2 Grounding Task
Comparing Figure 19 and Figure 19, it can be observed that Qwen-VL performs well on the original object detection tasks. However, after undergoing Stage 1 training, the LLM experienced a decline in performance on these tasks. However, comparing Figures 19 and 19, it can be seen that the model trained with Stage 1 shows improvement on the Mobile3M Grounding task compared to the original Qwen-VL. In the selected examples, Qwen had significant errors, but Stage 1 provided a corrective effect across a broader range of examples. Considering that the data is suitable for training specialized models in the Mobile domain, the performance loss in the general domain is acceptable.
@@ D.2 Grounding Task
D.3 Pre-training Task Test Case
@@
3. Action Space Generation As shown in Figure 21, based on the Element List, we expect the model to further distinguish the type of each widget. Since scroll elements are not visible themselves, we manually generated four alternative actions for them: up, down, left, and right. Our action space does not support diagonal scrolling, as this often implies a drag action in practical operations. Therefore, the model actually only needs to distinguish between input fields and buttons. We also expect it to ground these buttons accurately.
D.4 Fine-tuning Task Case Study
The limitation of the MobileVLM-unified model in the Auto-UI single task: Multi-step mixed training tasks cause the model to make more errors in determining the end of a task. Below are the different responses from MobileVLM-unified and MobileVLM-separate when completing the same task: